# Energy landscape interpretation of universal linearly increasing absorption with frequency

### v2

@@ IV.1 Validity of the Arrhenius relation
One thing is that properties are described by an energy landscape, another is the validity of the Arrhenius relation. As stated in [35], the Arrhenius relation comes from statistical mechanics and is valid for a system which transitions from one metastable state to another. They argue that although these assumptions are not necessarily valid in biological tissue, analogs to the relevant parameters “exist in cells and likely govern cell motility.”
Likewise, [36] argues that proteins also can be regarded as two-state systems. They cite as examples “an ion channel can be open or closed, a hemoglobin or myoglobin protein can have bound oxygen or not.” The Arrhenius law is therefore used in these fields as well, although the physical basis is not as solid as for processes at the atomic and molecular scale.
IV.2 Soft glassy materials and glass
@@
There is independent evidence that human tissue cells under the influence of shear may be modeled with soft glassy rheology [37], [15], [38]. This lends support to the just mentioned descriptions of biological materials,
@@
The properties of soft glasses correspond to those of glasses at temperatures between the glass temperature and the melting point, $T_{g}<T<T_{m}$. The cell’s mechanical properties are determined by the crowded interior of the cell. This is analogous to what takes place in a colloidal suspension and leads to the complex shear modulus following a weak power law over several frequency decades with a near constant power law exponent, similar to that of Supplement Eq. (S7). Cells are examples of a disordered metastable material which exists in a state far from thermodynamic equilibrium. The energy landscape is comprised of the binding energy between neighboring proteins. A deformation due to an incoming wave may cause a hop between energy wells where the deformation energy is taken from the wave, i.e. leading to heating and absorption of wave energy.
IV.3 Sensitivity to energy distribution in the band-limited case
@@
A high-frequency limit like that of (9) may in fact be required for physical reasons as a passive medium requires that the absorption should not increase faster than $\omega^{1}$ as $\omega$ approaches infinity, [40]. As noted, the low-frequency limit is relevant for modeling e.g., sub-bottom sediments, [41]. The exponential relationship of the Arrhenius law also means that a band-limiting corresponds to a relatively narrow range of activation energies.
@@ V Discussion
Glasses have a universal property at low temperatures, 0.1 to 10 K, where $Q^{-1}(\omega;T)$ is found to be nearly independent of temperature T as well as frequency $\omega$, [42]. The property that elastic wave absorption depends on frequency in a linear way around room temperature is a similar universal property which results from a flat activation energy distribution.
An alternative way of deriving an energy distribution similar to (8) is found in the theory for soft glassy materials, [43], [44] where the medium was modeled as a Maxwell-Wiechert model as in Fig. (S1) and Eq. (S5) of the Supplement.
@@
That theory also gives a material description in the form of a partial differential equation which expresses how regions rearrange to new positions, valid for low frequencies. The main variable is $g_{E}(E_{a})$, the probability for finding an element trapped in a barrier between the two wells of height $E_{a}$.
@@
The theory contains a constant which is an attempt frequency, and an activation factor on the same form as the Arrhenius equation. The theory is expressed in normalized units with a central parameter being the mean-field noise temperature, $y$. A glass transition occurs at $y=y_{g}=1$ and the material approaches the fluid state for $y=2$. Another input is a prior distribution of traps which it is argued has an exponential tail, $p_{a}(E_{a})=\exp(-E_{a}/y_{g})$, [45].
The equilibrium distribution of energies above the glass transition is given by $g_{E}(E_{a})\propto\mathrm{e}^{E_{a}/y}p_{a}(E_{a})$, and although the end result is not stated explicitly in [44, Sect. IV.A], it is an exponential distribution:
$g_{E}(E_{a})\propto\mathrm{e}^{\frac{-E_{a}}{y}(y-1)}.$ (11)
Since $y$ corresponds to $k_{b}T$, this expression is analogous to (8). In addition, it states that the parameter $y$ changes with temperature. New results are now appearing with ultrasound properties of tissue at low temperature that potentially could confirm this, [46].
@@
Further it is demonstrated how this model leads to a dynamic modulus ${E}(\omega)={E}^{\prime}(\omega)+\mathrm{i}{E}^{\prime\prime}(\omega)$ where both the real and the imaginary components are proportional to $\omega^{y-1}$, as for the fractional Kelvin-Voigt model of Supplement Fig. (S2) and Eq. (S7).
@@
Thus the noise temperature and the fractional order have a simple relationship, $y=\beta+1$. The soft glassy model therefore provides an interpretation of the fractional order, $\beta$ as well as for $y$. The special case of concern in this article is found in the limit, as $y$ approaches one, i.e. the material approaches the glass temperature. It is also evident that the soft glassy model result of (11) for the energy distribution resembles our result (8) which we found in an independent way.
As mentioned in the Introduction, the typical range for $y$ for shear waves in biological tissue is up to $1.5$. This corresponds to an energy distribution which is skewed towards lower energies, see Fig. 1. One explanation for why it is not flat could be that the Arrhenius relation is not completely valid for some biological materials. Variations such as the stretched exponential Arrhenius relation could be relevant for further investigation.
@@ VI Conclusion
Such absorption is the result of a large number of relaxation processes, expressed by a weighted relaxation integral over frequency or over time. It is however hard to argue physically why the particular weighting that gives rise to power-law absorption, should occur. We have argued here, based on properties of atomic and molecular clusters, proteins, and glasses that an energy landscape formulation is fundamental. Further the Arrhenius expression for activation energy, despite being formally derived in statistical mechanics, is valid. This enables the transformation of the multiple relaxation formulation to an integral over energies in an energy landscape. The formulation derived here is not common in e.g. ultrasonics and seismics.
A macroscopic property, absorption, is linked to properties at the mesoscale level, the shape of the energy landscape. The interesting case of linearly increasing absorption corresponds to a flat activation energy distribution, i.e., all energies are equally probable. A flat energy distribution indicates a form of equilibrium, and it is not unlikely that properties of the energy landscape will enable a deeper understanding of both the conditions for linearly increasing absorption with frequency as well as the origin of power-law relaxation responses in general.
VII Acknowledgments

### v3

@@ IV.1 Validity of the Arrhenius relation
One thing is that properties are described by an energy landscape, another is the validity of the Arrhenius relation for biological materials. As stated in [43], the Arrhenius relation comes from statistical mechanics and is valid for a system which transitions from one metastable state to another. They argue that although these assumptions are not necessarily valid in biological tissue, analogs to the relevant parameters “exist in cells and likely govern cell motility.”
The energy landscapes of proteins and glasses also have many properties in common. Proteins may in some cases be regarded as two-state systems, such as when “an ion channel can be open or closed, a hemoglobin or myoglobin protein can have bound oxygen or not” [44].
The Arrhenius law is therefore not uncommon to use for characterizing the energy landscapes of cells and proteins as well. Its validity is not as universal as in the fields discussed previously [45], and there are several unsolved questions in this field. We assume here that the Arrhenius equation may be applied, although the physical basis is not as solid as for processes at the atomic and molecular scale.
IV.2 Soft glassy materials
@@
There is independent evidence that human tissue cells under the influence of shear may be modeled with soft glassy rheology [46], [28], [47]. This lends support to the just mentioned descriptions of biological materials. The response often varies with $\omega^{\beta}$, where $\beta$ is in the range 0.1 to 0.5 [28], [48]. According to results cited in the Appendix, this corresponds to an absorption that varies with $y=\beta+1$ and therefore $y$ in this case will fall in the range 1.1 to 1.5.
@@
The properties of soft glasses correspond to those of glasses at temperatures between the glass temperature and the melting point, $T_{g}<T<T_{m}$. The cell’s mechanical properties are determined by the crowded interior of the cell. This is analogous to what takes place in a colloidal suspension and leads to the complex shear modulus following a weak power law over several frequency decades with a near constant power law exponent, similar to that of Appendix Eq. (25). Cells are very soft relative to the materials they are made from, similar to how a wool jumper is soft compared to the wool fibers that it comprises. Cells are examples of a disordered metastable material which exists in a state far from thermodynamic equilibrium. The energy landscape is comprised of the binding energy between neighboring proteins. An incoming wave may cause a hop from one of the states mentioned above to the other, where the required energy is taken from the wave, i.e. leading to heating and absorption of wave energy.
IV.3 Non-Arrhenius behavior
As mentioned in the previous section, the typical range for $y$ for shear waves in biological tissue is up to $1.5$. Although $y=1$ is the most common value for compressional waves in medical ultrasound, the value may reach up to $1.5$ even in this field [10]. If the Arrhenius equation is valid, this could mean that the energy distribution is skewed towards lower energies as in Fig. 1.
Alternatively, it could also mean that the Arrhenius equation is no longer valid for some biological materials or under specific conditions. The stretched exponential has been proposed as an alternative:
$\tau=\tau_{0}\mathrm{e}^{\left(E_{a}/k_{B}T\right)^{\gamma}},$ (14)
where the range $10<\gamma<$ 1 leads to a stretched exponential, $\gamma>1$ gives a compressed exponential, and $\gamma=2$ results in a gaussian distribution. Repeating the steps leading from (9) to (12), gives a new energy distribution:
$g_{E}(E_{a})=\frac{K_{y}{\tau_{0}^{1-y}}}{k_{B}T}\;\gamma\cdot\mathrm{e}^{-\left(\frac{E_{a}}{k_{B}T}\right)^{\gamma}(y-1)}\cdot\left(\frac{E_{a}}{k_{B}T}\right)^{\gamma-1},$ (15)
which for $\gamma=1$ contains (12) as a special case.
When $\gamma>1$, this equation may predict that the flattest energy distribution occurs for $y>1$, but the energy distribution is no longer exactly flat as is the case for $\gamma=1$ and $y=1$. As an example $\gamma=1.5$ results in the flattest energy distribution for $y=1.1$. Eq. (15) in combination with a flat energy landscape may therefore only partially explain power-law exponents, $y$, above 1.
IV.4 Band-limited power laws
@@
A high-frequency limit like that of (16), corresponding to a lower energy limit, may in fact be required for physical reasons as a passive medium requires that the absorption should not increase faster than $\omega^{1}$ as $\omega$ approaches infinity [50]. As noted, the low-frequency limit corresponding to an energy distribution which stops at a maximum energy value, is sometimes observed in e.g., sub-bottom sediments [51]. The exponential relationship of the Arrhenius law also means that a band-limiting corresponds to a relatively narrow range of activation energies.
@@ V Discussion
Glasses have a universal property at low temperatures, 0.1 to 10 K, where $Q^{-1}(\omega;T)$ is found to be nearly independent of temperature T as well as frequency $\omega$ [52]. The property that elastic wave absorption depends on frequency in a linear way around room temperature is a similar universal property. We have shown here that it results from a flat activation energy distribution.
@@
An alternative way of deriving an energy distribution similar to (12) is found in the theory for soft glassy materials [53], [54] where the medium was modeled as a Maxwell-Wiechert model as in Fig. 3 and Eq. (22) of the Appendix. That theory gives a material description in the form of a partial differential equation which expresses how regions rearrange to new positions, valid for low frequencies. The main variable is $g_{E}(E_{a})$, the probability for finding an element trapped in a barrier between the two wells of height $E_{a}$.
@@
The theory contains a constant which is an attempt frequency, and an activation factor on the same form as the Arrhenius equation. The theory is expressed in normalized units with a central parameter being the mean-field noise temperature, $y$. A glass transition occurs at $y=y_{g}=1$ and the material approaches the fluid state for $y=2$. Another input is a prior distribution of traps which it is argued has an exponential tail, $p_{a}(E_{a})=\exp(-E_{a}/y_{g})$ [55].
The equilibrium distribution of energies above the glass transition is given by $g_{E}(E_{a})\propto\mathrm{e}^{E_{a}/y}p_{a}(E_{a})$, and although the end result is not stated explicitly in [54, Sect. IV.A], it is an exponential distribution, $g_{E}(E_{a})\propto\mathrm{e}^{-E_{a}(y-1)/y}$. Since $y$ is $k_{b}T/E_{0}$ [40], this expression is analogous to (12). The dependence of $y$ on temperature outlines one way of testing (12) and new results with ultrasound properties of tissue at low temperature could potentially be used [56]. However the possible temperature range is rather limited compared to the range of temperature variation that glasses are tested under. The maximal temperature range that tissue can be subjected to without structural damage is 5–40 C [57].
@@
Further it is demonstrated how this model leads to a dynamic modulus ${E}(\omega)={E}^{\prime}(\omega)+\mathrm{i}{E}^{\prime\prime}(\omega)$ where both the real and the imaginary components are proportional to $\omega^{y-1}$, as for the fractional Kelvin-Voigt model of Appendix Fig. 4 and Eq. (25).
@@
Thus the noise temperature and the fractional order have a simple relationship, $y=\beta+1$. The soft glassy model therefore provides an interpretation of the fractional order, $\beta$ as well as for $y$. The special case of concern in this article is found in the limit, as $y$ approaches one, i.e. the material approaches the glass temperature. It is also evident that the soft glassy model result of [54] for the energy distribution resembles our result (12). In the soft glassy theory it is found in a bottom-up way, while our independent derivation is more of a top-down approach.
A philosophically inclined reader may have noted that we have not followed the often-desired hypothetico-deductive way of arguing strictly, as so far, only the temperature dependence of $y$ has been proposed as a way of partially testing our hypothesis. The main argument here is inference to best explanation, or abductive reasoning, arguing that an Arrhenius-type relation and a formulation in terms of an energy landscape unite many of the materials that display linearly increasing absorption, even those where other processes than thermal activation at the molecular level take place. Additionally, a flat energy landscape with maximal randomness makes sense as it can be interpreted as a maximum entropy distribution of energies.
@@ VI Conclusion
Such absorption is the result of a large number of relaxation processes, expressed by a weighted relaxation integral over frequency or over time. It is however hard to argue physically why the particular weighting that gives rise to power-law absorption, should occur. We have argued here, based on properties of atomic and molecular clusters, proteins, and glasses that an energy landscape formulation is fundamental. Further the Arrhenius expression for activation energy, despite being formally derived in statistical mechanics, is often valid. This enables the transformation of the multiple relaxation formulation to an integral over energies in an energy landscape.
A macroscopic property, absorption, is linked to properties at the mesoscale level, the shape of the energy landscape. The interesting case of linearly increasing absorption corresponds to a flat activation energy distribution, i.e., all energies are equally probable. A flat energy distribution indicates a form of equilibrium, and properties of the energy landscape may enable a deeper understanding of both the conditions for linearly increasing absorption with frequency as well as the origin of power-law relaxation responses in general.
Acknowledgments
@@
Appendix
There are ways of modeling power-law absorption that are different from the multiple relaxation model. They are referred to in some of the references of the main text and therefore two alternatives are discussed here: A mechanical model consisting of an infinite network of springs and dampers (Maxwell-Wiechert model), and a fractional Kelvin-Voigt model.
VI.1 Mechanical interpretation
Figure 3: Maxwell-Wiechert model with $E$’s denoting modulus of elasticity and $\eta$’s viscosity.
The multiple relaxation models can be given a mechanical interpretation by considering the discrete Maxwell-Wiechert model consisting of parallel branches of springs and dampers in series, as shown in Fig. 3. Its dynamic modulus is:
${E}(\omega)=E_{e}+\sum_{n=1}^{N-1}\frac{E_{n}\;\mathrm{i}\omega\tau_{n}}{1+\mathrm{i}\omega\tau_{n}}$ (18)
where $\tau_{n}=\eta_{n}/E_{n}$. Note that in this appendix $E$ means elasticity modulus and not energy as in the main text. The spring represented by $E_{e}$, the equilibrium value, is required in order to make this a model for a solid. In the limit, it is a continuous model where $\hat{E}(\tau)$ is the relaxation spectrum or distribution of elastic moduli [58, Sect. 4.1.1], whether they are shear or bulk moduli:
${E}(\omega)=E_{e}+\int_{0}^{\infty}\frac{\hat{E}(\tau)\;\mathrm{i}\omega}{1+\mathrm{i}\omega\tau}\;\mathrm{d}\tau,$ (19)
There is a direct relation to the absorption of (4) and (6) which can be found from linearized conservation of momentum and energy. In that case the dispersion relation can be shown to be [2, Sect. 3.5]:
$k^{2}(\omega)=\rho_{0}\frac{\omega^{2}}{E(\omega)}$ (20)
Absorption is found from the imaginary part of the wavenumber $k(\omega)$:
$\alpha(\omega)=-\sqrt{\frac{\rho_{0}}{E_{e}}}\omega\;\Im\left[1+\int_{0}^{\infty}\frac{\mathrm{i}\omega\;\hat{E}(\tau)/E_{e}}{1+\mathrm{i}\omega\tau}\;\mathrm{d}\tau\right]^{-1/2}$ (21)
Assuming that $\hat{E}(\tau)\ll E_{e}$, i.e., all loss mechanisms are weak, the absorption is:
$\alpha(\omega)\approx\sqrt{\frac{\rho_{0}E_{e}}{2}}\;\omega^{2}\int_{0}^{\infty}\frac{\hat{E}(\tau)}{1+\omega^{2}\tau^{2}}\;\mathrm{d}\tau$ (22)
which is similar to (6) with $\hat{E}(\tau)\propto g_{\tau}(\tau)\tau$. The exact expression for $\hat{E}(\tau)$ in the power-law case is:
$\hat{E}(\tau)=\sqrt{\frac{2}{\rho_{0}E_{e}}}\;K_{y}\tau^{-y}.$ (23)
This description therefore gives insight into how the relaxation integrals of (4) and (6) can be realized in terms of elementary mechanical models, but it does not give us much insight into why $\hat{E}(\tau)$ is shaped in a particular way in the case of power-law absorption.
VI.2 Fractional Kelvin model
Figure 4: Fractional Kelvin-Voigt medium model with $E_{e}$ denoting modulus of elasticity and $\eta$ denoting pseudo-viscosity of order $\beta$.
Fractional viscoelasticity is an alternative way of expressing power-law absorption. Its advantage is that it requires a small number of parameters.
In particular the fractional Kelvin-Voigt mechanical model shown in Fig. 4 has been used for tissue discrimination in elastography, i.e. shear wave imaging in tissue [59], and has been recommended for general use in elastography [48]. The model consists of a fractional damper of order $\beta$ in parallel with a spring [60, Sect. 3.1], [2, Sect. 5.2], where the stress is given by
$\sigma(t)=E_{e}\varepsilon(t)+\eta\;\frac{\partial^{\beta}\varepsilon(t)}{\partial t^{\beta}},$ (24)
where $\varepsilon(t)$ is strain, $E_{e}$ is the spring’s modulus of elasticity, and $\eta$ is a pseudo-viscosity in units Pa/sβ. The dynamic modulus, the Fourier transform of the stress impulse response, is
${E}(\omega)=E_{e}+\eta\;(\mathrm{i}\omega)^{\beta}.$ (25)
Here both the real and the imaginary parts have components that are proportional to $\omega^{\beta}$. It is shown in [2, Sect. 5.6] that the wave equation in this case has a solution which gives rise to an attenuation which follows $\omega^{\beta+1}$ for low frequencies. Thus the power-law absorption of (1) for this model is:
$\alpha(\omega)=\alpha_{0}\omega^{y}=\alpha_{0}\omega^{\beta+1}.$ (26)
Author Declarations
Conflict of Interest Statement
The authors have no conflicts to disclose.
Data Availability
No data was collected for this research.
Ethics Approval
No ethics approval was required for this research.